OpenStudy (anonymous):

What does the graph of f(x) = 3x^2 – 2x + 1 look like? A. a parabola that opens upward B. a parabola that opens downward C. a v-shaped graph that opens upward D. a v-shaped graph that opens downward @Cosmichaotic @mrdoldum

OpenStudy (anonymous):

A. a parabola that opens upward

OpenStudy (anonymous):

f(x) = 3x^2 – 2x + 1 or y =3(x^2 -2x/3)+1 y =3(x^2 -2x/3 +1/9 -1/9)+1 y =3(x^2 -2x/3 +1/9 )-1/9+1 y= 3*(x-1/3)^2 +8/9 y-8/9=4*3/4*(x-1/3)^2 which is of the form Y= 4(3/4)X^2 where Y= y-8/9 and x-1/3=X

OpenStudy (nirmalnema):

sign of x^2 is +ve so the graph is upward opening and it is a quadratic function always parabola.. so its upward opening parabola

OpenStudy (anonymous):

Thank you both :)